One number is more than 3 times another. Their sum is 21. What is their product?
Answer by janinecb(25)

I think there is a number missing from this problem. I am going to guess which number is missing and then solve that problem for you. Hopefully this will help to solve the actual original problem.

The revised problem:
One number is 5 more than three times another. Their sum is 21. What is their product?

Turn the two first sentences into euqations. Here is the first one:
One number is 5 more than three times another.
x = y + 5

Notice that the word is turns into the equals sign and the more than turns into addition.

The second sentences is here:
Their sum is 21.
x + y = 21

Notice that the word sum signals addition and again the word is turns into the equals sign.

Once the two equations are set, use substitution to find one of the numbers.
(y + 5) + y = 21 (Note: y + 6 is being used instead of x in the second equation.)

This simplifies to:
2y + 5 = 21

The next step is to subtract 5 from both sides. This leaves you with:
2y = 16

Next divide both sides by 2. This leaves you with:
y = 8

Now that we found y, we can substitute 8 in for one of the y's in the first two equations to find out what x is.
x = y + 5 turns into x = 8 + 5

So, x = 13.

Now that we know that x = 13 and y = 8, we can answer the final question of what is the product of these two numbers.

Product means the answer to a multiplication problem, so we multiply 13 and 8 to get the final answer of 104.

Remember, this is with the assumption that the missing number in the problem is 5.